# Bosch China Investment Ltd

### v4

@@ 3.2.1 Order-Preserving Message Encoder
The main challenge in message encoding lies in effectively converting secret message bits into message features while retaining the sequential information of the message. To address this, we propose an OPME that efficiently encodes the message while preserving both local and global sequential information. Our model consists of three OPME in total to enable multi-scale interaction between the message and image features.
Each OPME incorporates an MLP, referred to as $\text{MLP}_{\text{enc}}$, which functions as a channel encoder. Its purpose is to encode message segments $M_{\text{seg}}\in\left\{0,1\right\}^{N_{\text{ms}}\times L_{\text{ms}}}$ into message feature $F_{\text{msg}}$ following Eq. 1:
@@
where the depth of $F_{\text{msg}}$ is greater than $L_{\text{ms}}$ and the shape of $F_{\text{msg}}$ is equal to the shape of $F_{\text{im}}$ at the corresponding GMIF. This learning-based channel encoder [O’shea and Hoydis(2017)] introduces redundant code into the original message segments, facilitating the stable transmission of the message feature within the model.
The importance of maintaining the sequential order of message segments is highlighted by the incorporation of PEs into the encoded segments. This is demonstrated in Eq. 2:
@@
where positional embedding $E_{\text{pos}}$ is introduced to the $F_{\text{msg}}$ followed by a MHSA module [Vaswani et al.(2017a)Vaswani, Shazeer, Parmar, Uszkoreit, Jones, Gomez, Kaiser, and Polosukhin] to globally encode $F_{\text{msg}}$ to $F_{\text{msg}}^{\text{global}}$. Layer Normalization (LN) [Ba et al.(2016)Ba, Kiros, and Hinton] is employed to normalize the global message features. We further validate effectiveness of PEs in Sec. 4.4.
@@ 3.2.2 Order-Preserving Message Decoder
As depicted in Fig. 2(b), we employ a standard SwinTransformer [Liu et al.(2021)Liu, Lin, Cao, Hu, Wei, Zhang, Lin, and Guo] with minor modifications to extract features from $I_{\text{stego}}$. Specifically, the down-sampling operations in the last two layers of the W-MSA module within SwinTransformer are removed to match the spatial size to the feature of the secret message in the OPME. Subsequently, the OPMD reconstructs the secret message segments $\hat{M}$ from the extracted features. The OPMD follows the same methodology as the OPME, but in reverse order.
@@ 3.3 Global Message-Image Fusion
is designed to accomplish two fundamental goals of message hiding. These goals encompass: (1) the stego image should contain all the information from both the secret message and the cover image; (2) the ideal stego image should be nearly identical to the cover image. The key, value and query in the attention mechanism are configured according to these two goals to translate the message and image information to the steganographic features where the key, value and query are set as: $K=(F_{\text{im}}+F_{\text{msg}}^{\text{global}})W^{K}$, $V=(F_{\text{im}}+F_{\text{msg}}^{\text{global}})W^{V}$, and $Q=F_{\text{im}}W^{Q}$. This configuration utilized in our approach are inspired by the Transformer-based architecture commonly employed in language translation tasks. Drawing a parallel, we apply a similar mechanism to the task of message hiding, as both involve dealing with a lack of spatial and semantic alignment between the secret message and cover image. The GMIFeffectively fuses the optimal stenographic information, which is the sum of $F_{\text{msg}}^{\text{global}}$ and $F_{\text{im}}$, into the desired feature representation of the stego image $F_{\text{stego}}$. More details related to the key, value, and query of GMIF are provided in the supplementary material.
The basic building block of GMIF  is W-MSA. One layer of W-MSA with large window size is applied to ensure a comprehensive global interaction between message and image features in the GMIF. We apply up-sampling and convolution layers to the concatenated $F_{\text{stego}}$ and $F_{\text{im}}$ to reconstruct the residual image $I_{\text{res}}$. The up-sampling layer uses a de-convolution operation to double the height and width of the feature map and reduce the feature dimension by half. The final output of our model is $I_{\text{stego}}$, which is equal to the sum of $I_{\text{cover}}$ and $I_{\text{res}}$. Please check supplementary material for detailed configurations of our model.
@@ 3.4 Loss Function
The objectives of message hiding are two-fold: to generate a stego image that closely resembles the cover image and to precisely retrieve the secret message from the stego image. In pursuit of these aims, we employ the loss function defined as $\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{img}}+\mathcal{L}_{\text{msg}}$, where $\mathcal{L}_{\text{img}}=\lambda_{1}\mathcal{L}_{\text{MSE}}(I_{\text{cover}},I_{\text{stego}})+\lambda_{2}\mathcal{L}_{\text{LPIPs}}(I_{\text{cover}},I_{\text{stego}})$ where LPIPs [Zhang et al.(2018)Zhang, Isola, Efros, Shechtman, and Wang] is perceptual loss, and $\mathcal{L}_{\text{msg}}=\mathcal{L}_{\text{BCE}}(M,\hat{M})$ if the range of message $N_{r}=1$, or $\mathcal{L}_{\text{msg}}=\mathcal{L}_{\text{MSE}}(M,\hat{M})$ if the message range $N_{r}>1$.
Dataset COCO
Metric ACC PSNR SSIM Inference Time
Capacity 1 BPP 2 BPP 3 BPP 1 BPP 2 BPP 3 BPP 1 BPP 2 BPP 3 BPP ms
LISO* [Chen et al.(2023)Chen, Kishore, and Weinberger] 99.99% 99.98% 99.84% 33.67 33.40 28.20 0.9424 0.9350 0.8184 607.7
SteganoGAN [Zhang et al.(2019a)Zhang, Cuesta-Infante, and Veeramachaneni] 97.75% 96.46% 91.35% 42.52 39.68 36.45 0.9894 0.9801 0.9650 1.6
ChatGAN [Tan et al.(2021)Tan, Liao, Liu, Cao, and Jiang] 99.07% 97.46% 96.18% 46.42 43.17 41.84 0.9943 0.9880 0.9832 3.0
StegaFormer (Ours) 99.95% 99.85% 99.68% 47.83 45.30 43.37 0.9969 0.9943 0.9914 9.1
Dataset DIV2K
Metric ACC PSNR SSIM Inference Time
Capacity 1 BPP 2 BPP 3 BPP 1 BPP 2 BPP 3 BPP 1 BPP 2 BPP 3 BPP ms
LISO* [Chen et al.(2023)Chen, Kishore, and Weinberger] 99.91% 99.83% 99.55% 35.52 33.67 30.13 0.9394 0.9088 0.8326 607.7
SteganoGAN [Zhang et al.(2019a)Zhang, Cuesta-Infante, and Veeramachaneni] 98.96% 97.63% 92.23% 39.36 37.97 36.14 0.9793 0.8916 0.8343 1.6
ChatGAN [Tan et al.(2021)Tan, Liao, Liu, Cao, and Jiang] 99.73% 98.63% 94.75% 45.00 42.05 40.63 0.9929 0.9861 0.9782 3.0
StegaFormer (Ours) 99.94% 99.87% 99.76% 51.98 49.07 47.31 0.9985 0.9970 0.9958 9.1
Table 1: Quantitative comparisons of different approaches with ours on the COCO and DIV2K datasets are presented. ACC denotes message recovery accuracy, while PSNR and SSIM are two metrics to measure the level of imperceptibility. The best results for non-iterative methods are highlighted in bold. * denotes iterative method.
@@ 4 Experiments
To demonstrate the effectiveness of our proposed approach, we conduct a comparative analysis of message recovery accuracy and imperceptibility against three state-of-the-art message hiding methods. SteganoGAN [Zhang et al.(2019a)Zhang, Cuesta-Infante, and Veeramachaneni], ChatGAN [Tan et al.(2021)Tan, Liao, Liu, Cao, and Jiang] are non-iterative approaches and LISO [Chen et al.(2023)Chen, Kishore, and Weinberger] is iterative. This analysis is carried out across different message capacity levels.
Following previous studies [Zhang et al.(2019a)Zhang, Cuesta-Infante, and Veeramachaneni, Tan et al.(2021)Tan, Liao, Liu, Cao, and Jiang, Chen et al.(2023)Chen, Kishore, and Weinberger], our method is evaluated on two public datasets: COCO [Lin et al.(2014)Lin, Maire, Belongie, Hays, Perona, Ramanan, Dollár, and Zitnick] and DIV2K [Agustsson and Timofte(2017)]. The COCO dataset for message hiding contains 25,000 images randomly sampled from the original COCO training split for model training, as well as 500 images randomly sampled from the original COCO test split for model evaluation. No extra configurations for DIV2K dataset. All images are center cropped to $256\times 256$ pixels. Three metrics are used to evaluate the model performance. The message recovery accuracy is utilized to assess the percentage of successfully recovered secret messages. The imperceptibility of stego images is measured using the PSNR and SSIM. More detailed configurations of our model are listed in the supplementary material due to the page limitation.
@@ 4.1 Comparison with State-of-the-Art Methods
We conducted a comprehensive comparison between our method and two existing non-iterative message hiding methods, namely SteganoGAN [Zhang et al.(2019a)Zhang, Cuesta-Infante, and Veeramachaneni] and ChatGAN [Tan et al.(2021)Tan, Liao, Liu, Cao, and Jiang], as well as the latest iterative method LISO [Chen et al.(2023)Chen, Kishore, and Weinberger]. As summarized in Tab. 1, our method consistently outperforms the non-iterative methods in terms of recovery accuracy and imperceptibility across different message capacities (1, 2, and 3 BPP). Specifically, at a 3 BPP capacity, our method achieves impressive recovery accuracies of $99.68\%$ and $99.76\%$ for the COCO and DIV2K datasets, respectively. These results represent significant improvements of $3.5\%$ and $5.0\%$ over the current non-iterative state-of-the-art accuracies achieved by ChatGAN, demonstrating the effectiveness of our approach (OPME) in achieving high-accuracy message hiding. In contrast to the iterative method, our approach achieves comparable levels of message recovery accuracy. Notably, our method outperforms the accuracy of iterative approach when trained on high quality image datasets such as DIV2K.
Furthermore, our method enhances imperceptibility in terms of both PSNR and SSIM across different message capacities. It achieve much higher quality than iterative method since this approach significantly hinder the quality of stego image. Compared with the best non-iterative method ChatGAN, at a 3 BPP message capacity, our method improves the PSNR from 41.84 to 43.37 on the COCO dataset and from 40.63 to 47.31 on the DIV2K dataset. These results underscore the effectiveness of our approach (GMIF) in fusing steganographic information into stego images, achieving significantly higher imperceptibility than the iterative method. In terms of inference time, our method achieves similar levels of efficiency as other non-iterative approaches. In contrast, iterative methods such as LISO typically require over 600 ms due to the extra iterations during inference.
Fig. 3 illustrates a cover image (a), stego images (b) to (d) generated by our method and previous methods [Zhang et al.(2019a)Zhang, Cuesta-Infante, and Veeramachaneni, Tan et al.(2021)Tan, Liao, Liu, Cao, and Jiang] at 3 BPP message capacity, and residual images (e) to (g). All the pixel values of residual images are enlarged by 5 times. The residual images produced by our method demonstrate that minimal changes have been applied to the cover images. In contrast, previous approaches introduce higher residuals throughout the cover image, resulting in lower imperceptibility. More qualitative results are in the supplementary material.
$I_{cover}$
(a) $I_{stego}$
(b) $I_{stego}$
(c) $I_{stego}$
(a) $I_{res}$
(b) $I_{res}$
(c) $I_{res}$
Figure 3: Qualitative results of 3 BPP message hiding using different methods: (a) SteganoGAN, (b) ChatGAN and (c) Ours. $I_{cover}$, $I_{stego}$ and $I_{res}$ reperesent cover image, stego image and residual image respectively.
$I_{cover}$
(a) $I_{stego}$
(b) $I_{stego}$
(c) $I_{stego}$
(a) $I_{res}$
(b) $I_{res}$
(c) $I_{res}$
Figure 4: Qualitative results of high-capacity message hiding with our method: (a) 4 BPP, (b) 6 BPP and (c) 8 BPP. $I_{cover}$, $I_{stego}$ and $I_{res}$ represent cover image, stego image and residual image respectively.
@@ 4.2 Models for High Message Capacity
We conduct a series of experiments to showcase the performance of our method in handling high message capacities. As shown in Tab. 3, the results demonstrate that our method achieves more than $90\%$ message recovery accuracy for message capacities up to 8 BPP. Compared with existing methods, these results highlight the remarkable improvements of our method in both message capacity and recovery accuracy. Besides, Fig. 4 (b) to (g) depicts the stego and residual images by our models with high message capacities. The residual images of these high-capacity models demonstrate the outstanding imperceptibility of our models. Please check the supplementary material for details related to large message capacity models.
4.3 Steganalysis
Steganalysis methods [Boroumand et al.(2018)Boroumand, Chen, and Fridrich, Zhang et al.(2019b)Zhang, Zhu, Liu, and Liu, You et al.(2020)You, Zhang, and Zhao] play a crucial role in assessing the security of stego images. In our study, we choose the state-of-the-art steganalysis approach SiaStegNet [You et al.(2020)You, Zhang, and Zhao] to conduct a comprehensive benchmark comparison between our method and existing methods [Zhang et al.(2019a)Zhang, Cuesta-Infante, and Veeramachaneni, Tan et al.(2021)Tan, Liao, Liu, Cao, and Jiang, Chen et al.(2023)Chen, Kishore, and Weinberger]. Fig. 6 presents the steganalysis results on COCO dataset, demonstrating that our model achieves the lowest Area Under the Curve (AUC) among the three compared methods. These results indicates that our method exhibits the lowest detection rate and has the best security of stego images. All message hiding models for comparison have a 3 BPP message capacity. More details of steganalysis are in the supplementary material.
Figure 5: Detection accuracy of different methods. Lower is better.
Figure 6: The eigen values analysis of PEs from ViT and OPME.
@@ 4.3 Steganalysis
Table 2: StegaFormer results at large BPPs.
Model ACC PSNR SSIM
Baseline 50.01 53.4 0.9988
+GMIF 87.60 43.78 0.9908
+GMIF+OPME+OPMD 99.24 41.78 0.9875
Table 3: Ablation of OPME, OPMD and GMIF.

### v5

@@ 3.2.1 Order-Preserving Message Encoder
The main challenge in message encoding lies in effectively converting secret message bits into message features while retaining the sequential information of the message. To address this, we propose an OPME that efficiently encodes the message while preserving both local and global sequential information. In this context, “local" refers to the order within a message segment, while “global" refers to the sequential order of the message segments. Our model consists of three OPME in total to enable multi-scale interaction between the message and image features.
Each OPME incorporates an MLP, referred to as $\text{MLP}_{\text{enc}}$, which functions as a channel encoder. Its purpose is to encode message segments $M_{\text{seg}}\in\left\{0,1\right\}^{N_{\text{ms}}\times L_{\text{ms}}}$ into message feature $F_{\text{msg}}$ following Eq.˜1:
@@
where the shape of $F_{\text{msg}}$ is equal to the shape of $F_{\text{im}}$ at the corresponding GMIF. This learning-based channel encoder O’shea and Hoydis (2017) introduces redundancy into the original message segments, facilitating the stable transmission of the message feature within the model.
The importance of maintaining the sequential order of message segments is highlighted by the incorporation of PEs into the encoded segments. This is demonstrated in Eq.˜2:
@@
where positional embedding $E_{\text{pos}}$ is introduced to the $F_{\text{msg}}$ followed by a MSA module Vaswani et al. (2017a) to globally encode $F_{\text{msg}}$ to $F_{\text{msg}}^{\text{global}}$. Layer Normalization (LN) Ba et al. (2016) is employed to normalize the global message features.
@@ 3.2.2 Order-Preserving Message Decoder
As depicted in Fig.˜3(b), we employ a standard SwinTransformer Liu et al. (2021) with minor modifications to extract features from $I_{\text{stego}}$. Specifically, the down-sampling operations in the last two layers of the W-MSA module within SwinTransformer are removed to match the spatial size to the feature of the secret message in the OPME. Subsequently, the OPMD reconstructs the secret message segments $\hat{M}$ from the extracted features. The OPMD follows the same methodology as the OPME, but in reverse order.
@@ 3.3 Global Message-Image Fusion
is designed to accomplish two fundamental goals of message hiding. These goals encompass: (1) the stego image should contain all the information from both the secret message and the cover image; (2) the ideal stego image should be identical to the cover image. The key, value and query in the attention mechanism are configured according to these two goals to translate the message and image information to the steganographic features $F_{\text{stego}}$:
$K$ $=(F_{\text{im}}+F_{\text{msg}}^{\text{global}})W^{K},$ (4)
$V$ $=(F_{\text{im}}+F_{\text{msg}}^{\text{global}})W^{V},$ (5)
$Q$ $=F_{\text{im}}W^{Q}.$ (6)
The configuration of query, key and value utilized in our approach are inspired by the Transformer-based architecture commonly employed in language translation tasks. Drawing a parallel, we apply a similar mechanism to the task of message hiding sharing, as both involve dealing with a lack of spatial and semantic alignment between the secret message and cover image. The GMIF effectively fuses the optimal stenographic information, which is the sum of $F_{\text{msg}}^{\text{global}}$ and $F_{\text{im}}$, into the desired feature representation of the stego image $F_{\text{stego}}$. Further insights into the various combinations of key, value, and query are provided in the supplementary material.
The basic building block of GMIF is Windowed Multi-Head Self-Attention Layers (W-MSA) Liu et al. (2021). One layer of W-MSA with large window size (all window sizes equal to 16) is applied to ensure a comprehensive global interaction between message and image features in the GMIF.
As shown in Eq.˜7, we apply up-sampling and convolution layers to the concatenated $F_{\text{stego}}$ and $F_{\text{im}}$:
$I_{\text{res}}=\text{Conv}(\text{Up-Sample}(\text{{Concat}}(F_{\text{stego}},F_{\text{im}}))),$ (7)
to reconstruct the residual image $I_{\text{res}}$. The up-sampling layer uses a de-convolution operation to double the height and width of the feature map and reduce the feature dimension by half. The final output of our model is $I_{\text{stego}}$, which is equal to the sum of $I_{\text{cover}}$ and $I_{\text{res}}$. Please check supplementary material for more detailed configurations and number of parameters of our model.
@@ 3.4 Loss Function
The objectives of message hiding are two-fold: to generate a stego image that closely resembles the cover image and to precisely retrieve the secret message from the stego image. In pursuit of these aims, we employ the loss function as defined in Eq.˜8:
$\mathcal{L}_{\text{total}}$ $=\mathcal{L}_{\text{img}}+\mathcal{L}_{\text{msg}},$ (8)
$\mathcal{L}_{\text{img}}$ $=\lambda_{1}\mathcal{L}_{\text{MSE}}(I_{\text{cover}},I_{\text{stego}})+\lambda_{2}\mathcal{L}_{\text{LPIPs}}(I_{\text{cover}},I_{\text{stego}}),$ (9)
$\mathcal{L}_{\text{msg}}$ $=\begin{cases}\mathcal{L}_{\text{BCE}}(M,\hat{M}),&\text{if}~N_{r}=1\\ \mathcal{L}_{\text{MSE}}(M,\hat{M}).&\text{if}~N_{r}>1\end{cases}$ (10)
$\mathcal{L}_{\text{img}}$ in Eq.˜9 combines a weighted sum of mean square error loss and perceptual loss Zhang et al. (2018). We employ binary cross-entropy loss for $\mathcal{L}_{\text{msg}}$ in Eq.˜10 when the range of message $N_{r}=1$ and $\mathcal{L}_{\text{msg}}$ change to mean square error loss when the message range $N_{r}>1$.
@@ 4 Experiments
To demonstrate the effectiveness of our proposed approach, we conduct a comparative analysis of message recovery accuracy and imperceptibility against two state-of-the-art message hiding methods: SteganoGAN Zhang et al. (2019a) and ChatGAN Tan et al. (2021). This analysis is carried out across different message capacity levels.
Experimental Settings. Following previous studies Zhang et al. (2019a); Tan et al. (2021), our method is evaluated on two public datasets: COCO Lin et al. (2014) and DIV2K Agustsson and Timofte (2017). We chose the COCO dataset due to its inclusion of diverse scenes and rich contextual image information, while the DIV2K dataset consists of high-resolution images. The COCO dataset for message hiding contains 25,000 images randomly sampled from the original COCO training split for model training, as well as 500 images randomly sampled from the original COCO test split for model evaluation. The DIV2K dataset contains 800 training images and 100 evaluation images. Throughout both the training and evaluation phases, we fix the image size to $256\times 256$ pixels for both datasets. The training process for each model consists of 100,000 iterations, with a batch size of 2. We set $\lambda_{1}=1\times 10^{-4}$ and $\lambda_{2}=1\times 10^{-6}$ to balance these different losses. More detailed configurations of our model are listed in the supplementary material due to the page limitation.
Evaluation Metrics. Three metrics are used to evaluate the model performance. The message recovery accuracy is utilized to assess the percentage of successfully recovered secret messages. The imperceptibility of stego images is measured using the PSNR and SSIM Wang et al. (2004).
@@ 4.1 Comparison with State-of-the-Art Methods
A comprehensive comparison between our method and two existing state-of-the-art methods Zhang et al. (2019a); Tan et al. (2021) is presented in Tab.˜1. Our method consistently outperforms the other methods in terms of recovery accuracy and imperceptibility at 1,2,3 BPP message capacity.
Specifically, at a 3 BPP capacity, our method achieves an impressive recovery accuracy of $99.68\%$ and $99.76\%$ for the COCO and DIV2K datasets, respectively. This represents a significant improvement of $3.5\%$ and $5.0\%$ compared to the current state-of-the-art accuracies achieved by ChatGAN Tan et al. (2021). These improvements demonstrate the effectiveness of OPME in achieving high-accuracy message hiding.
Furthermore, our method enhances imperceptibility in terms of both PSNR and SSIM across different message capacities. For example, compared with ChatGAN Tan et al. (2021) at a 3 BPP message capacity, our method improves the PSNR from 41.84 to 43.37 on the COCO dataset, and from 40.63 to 47.31 on the DIV2K dataset. These results underscore the effectiveness of GMIF in fusion of steganographic information into stego images.
Fig.˜5(b) displays residual images generated by our method and previous methods Zhang et al. (2019a); Tan et al. (2021) at 3 BPP message capacity. Residual images of our method demonstrate that minimal changes have been added to cover images. In contrast, previous approaches spread higher residuals all over the cover image, leading to lower imperceptibility. More qualitative results are in the supplementary material.
@@ 4.2 Models for High Message Capacity
Table 2: Quantitative results of StegaFormer with message capacities at 4, 6 and 8 BPP on the COCO dataset.
Figure 4: Qualitative results of 3 bits per pixel (BPP) message hiding using different methods: (a) cover image, (b) residual images, and (c) stego images. The images in the top row are obtained from SteganoGAN, those in the middle row from ChatGAN, and the bottom row showcases our method. For enhanced visualization, the residual images are multiplied by 100. Best viewed in digital version.
Figure 5: Qualitative results of high-capacity message hiding with our method: (a) cover image, (b) residual images, and (c) stego images. Samples at 4 BPP, 6 BPP, and 8 BPP message capacities are displayed in the first, second, and third rows, respectively. The residual images are multiplied by 100 for visualization. Best viewed in digital version.
We conduct a series of experiments to showcase the performance of our method in handling high message capacities. As shown in Tab.˜2, the results demonstrate that our method achieves a message recovery accuracy exceeding $90\%$ for message capacities up to 8 BPP. Compared with existing methods, these results highlight the remarkable improvement achieved by our method in both message capacity and recovery accuracy. Besides, Fig.˜5(b)(c) depicts the residual and stego images by our models with high message capacities. The residual images of these high-capacity models demonstrate the outstanding imperceptibility of our models.
There are two approaches to increase the message hiding capacity in our method: increasing the length of message segments $L_{\text{ms}}$ and expanding the range of message elements $N_{r}$. The 4 BPP model has $(L_{\text{ms}},N_{r})=(64,1)$, the 6 BPP model has $(L_{\text{s}},N_{r})=(48,3)$, and the 8 BPP model has $(L_{\text{ms}},N_{r})=(32,15)$. More details of $N_{r}$ and $L_{\text{ms}}$ are provided in the supplementary material.
4.3 Steganalysis
Steganalysis methods Boroumand et al. (2018); Zhang et al. (2019b); You et al. (2020) play a crucial role in assessing the security of stego images. Typically, this involves utilizing a machine learning model to classify images as either cover images or stego images. In our study, we choose the state-of-the-art steganalysis approach SiaStegNet You et al. (2020) to conduct a comprehensive benchmark comparison between our method and existing methods Zhang et al. (2019a); Tan et al. (2021). The SiaStegNet models are trained and evaluated on the COCO dataset.
Fig.˜6 presents the steganalysis results, demonstrating that our model achieves the lowest Area Under the Curve (AUC) among the three compared methods. These results indicates that our method exhibits the lowest detection rate and has the best security of stego images. All message hiding models for comparison have a 3 BPP message capacity.
Figure 6: The comparison of the detection accuracy of SiaStegNet models for ChatGAN, SteganoGAN and our approach. Lower AUC represents the stego images by our StegaFormer are more difficult to detect than the other two methods. Our approach shows lowest AUC.
Figure 7: The comparison of the ratio between the cumulative sum of the top n eigenvalues and the sum of all eigenvalues. Green for the PEs from pre-trained ViT and red for the PEs from OPME. The PEs from OPME contain more independent positional information.